# A unified description of curvature, torsion, and non-metricity of the metric-affine geometry with the Möbius representation

### v1

@@ Abstract
We establish the mathematical fundamentals for a unified description of curvature, torsion, and non-metricity 2-forms in the way extending the so-called Möbius representation of the affine group, which is the method to convert the semi-direct product into the ordinary matrix product, to revive the fertility of gauge theories of gravity. First of all, a bundle theory, focusing in particular on the Ehresmann connection and its curvature, is provided in a self-contained manner. Based on the theory, curvature, torsion, and non-metricity are introduced in the ordinal manner. Then the curvature and torsion 2-forms are described in a unified manner by using the Cartan connection of the Möbius representation of the affine group. After that, extending the Möbius representation, the dilation and shear 2-forms, or equivalently, the non-metricity 2-form, are introduced in the same unified manner. Finally, based on the unified description established in this paper, applying the Inönü-Wigner group contraction, the relationships among geometric quantities, gauge structures, and geometries are investigated with respect to the three gauge groups: the metric-affine group and its extension, and an extension of the (anti)-de Sitter group in which the non-metricity exists.
Keywords: gauge theories of gravity, metric-affine geometry, curvature, torsion, non-metricity, Ehresmann connection, Möbius representation, Inönü-Wigner contraction
@@ 1 Introduction
General Relativity (GR), which is, on one hand, the most successful theory of gravity, employs the (pseudo-)Riemannian geometry to describe gravitational phenomena based on Einstein’s equivalence principle and the general covariance [1, 2]. On the other hand, Yang-Mills gauge field theory, which describes fundamental interactions in the elementary particle physics such as electromagnetic, weak, and strong interactions in a unified manner based on the gauge principle, makes use of a principal bundle with a specific gauge group for each theory [3, 4, 5, 6, 7, 8]. Inspired by these studies, Utiyama tried to establish a unified description of all the above phenomena that is disciplined by the gauge invariant characteristics [9]. This is the dawn of the gauge theories of gravity. A detailed and historical development of this approach is illustrated in Refs [10, 11].
@@
Metric-Affine Gauge theories of gravity (MAG) are one of such approach to gauge theories of gravity [12, 13, 14]. This theory is constructed upon the metric-affine geometry of which geometric quantities are not only curvature but also torsion and non-metricity (the latter one is further decomposed into dilation and shear). Therefore, MAG is an extension/modification of GR in terms of the gauge theories. In particular, as a modification of GR, a set of subclasses of MAG labeled as the Geometrical Trinity of Gravity (GTG) has been vigorously investigated by virtue of its equivalence in the equations of motion to GR up to boundary terms in the meaning of the action functional [15]. Under the imposition of the so-called Weitzenböch gauge condition [16, 17, 18, 19, 20, 21], which implies that the curvature vanishes and then the geometry is flat, GTG has two subclasses: Symmetric Teleparallel Equivalent to GR (STEGR) and Teleparallel Equivalent to GR (TEGR).
@@
STEGR is, on one hand, obtained by the imposition of vanishing torsion that is realized in the use of the so-called Stükelberg fields [15]. This theory describes gravity in terms only of the non-metricity. Furthermore, the imposition of the so-called coincident gauge leads to a specific subclass that is equivalent to GR with excepting the second-order derivative terms of its action functional [22, 20]. It is nothing but what Einstein first treated for deriving his field equations when applying the variational principle [23], although it implies also that this subclass of GTG is deprived of the property of the covariance [24]. On the other hand, TEGR is obtained by the imposition of vanishing non-metricity and describes gravity in terms only of the torsion. However, differing from STEGR, unveiling the true identity of vanishing non-metricity from the viewpoint of its gauge structure is still under development. One of the reasons for this situation would be the lack of a unified description of non-metricity with curvature and torsion. In fact, geometries that are characterized by both curvature and torsion have a unified description labeled as Riemann-Cartan geometry [25, 26, 27]. In this geometry, these geometric quantities are constructed by the so-called Cartan connection, which is the extension of the Ehresmann connection by using the Möbius representation [28, 29, 14], of a principal bundle with the affine gauge group. This approach to constructing the geometries makes it possible to unveil its gauge structure and the true identity of vanishing torsion (and also vanishing curvature). Therefore, a unified description in the same manner that takes also non-metricity into account would open the way for unveiling the true identity of vanishing non-metricity. In addition, once this approach is realized, the unified description would make it possible to scrutinize the gauge structure not only of GTG but also of MAG. Furthermore, just alternating the gauge group, the unified description allows us to easily switch the metric-affine geometry to another geometry. Combining the so-called Inönü-Wigner group contraction [30, 31], the relationships among each geometry could be unveiled, and a new geometry, or equivalently, a new geometric quantity, would also be discovered. The purpose of the current paper is to provide the mathematical fundamentals of a unified description of curvature, torsion, and non-metricity for the above investigations.

### v2

# A unified-description of curvature, torsion, and non-metricity of the metric-affine geometry with the Möbius representation

@@ Abstract
We establish the mathematical fundamentals for a unified description of curvature, torsion, and non-metricity 2-forms in the way extending the so-called Möbius representation of the affine group, which is the method to convert the semi-direct product into the ordinary matrix product, to revive the fertility of gauge theories of gravity. First of all, we illustrate the basic concepts for constructing the metric-affine geometry. Then the curvature and torsion 2-forms are described in a unified manner by using the Cartan connection of the Möbius representation of the affine group. In this unified-description, the curvature and torsion are derived by Cartan’s structure equation with respect to a common connection 1-form. After that, extending the Möbius representation, the dilation and shear 2-forms, or equivalently, the non-metricity 2-form, are introduced in the same unified manner. Based on the unified-description established in this paper, introducing a new group parametrization and applying the Inönü-Wigner group contraction to the full theory, the relationships among symmetries, geometric quantities, and geometries are investigated with respect to the three gauge groups: the metric-affine group and its extension, and an extension of the (anti)-de Sitter group in which the non-metricity exists. Finally, possible applications to theories of gravity are briefly discussed.
Keywords: gauge theories of gravity, metric-affine geometry, curvature, torsion, non-metricity, Möbius representation, Inönü-Wigner contraction
@@ 1 Introduction
General Relativity (GR), which is, on one hand, the most successful theory of gravity, employs the (pseudo-)Riemannian geometry to describe gravitational phenomena based on Einstein’s equivalence principle and the general covariance [1, 2]. On the other hand, Yang-Mills gauge field theory, which describes fundamental interactions in the elementary particle physics such as electromagnetic, weak, and strong interactions in a unified manner based on the gauge principle, makes use of a principal bundle with a specific gauge group for each theory [3, 4, 5, 6, 7, 8]. Inspired by these studies, Utiyama tried to establish a unified-description of all the above phenomena that is disciplined by the gauge invariant characteristics. [9] This is the dawn of the gauge theories of gravity. A detailed and historical development of this approach is illustrated in Refs [10, 11].
@@
Metric-Affine gauge theories of Gravity (MAG) are one of such approach to gauge theories of gravity [12, 13, 14]. This theory is constructed upon the metric-affine geometry of which geometric quantities are not only curvature but also torsion and non-metricity (the latter one is further decomposed into dilation and shear). Therefore, MAG is an extension/modification of GR in terms of the gauge theories. In particular, as a modification of GR, a set of subclasses of MAG labeled as the Geometrical Trinity of Gravity (GTG) has been vigorously investigated by virtue of its equivalence in the equations of motion to GR up to boundary terms in the meaning of the action functional [15]. Under the imposition of the so-called Weitzenböch gauge condition [16, 17, 18, 19, 20], which implies that the curvature vanishes and then the geometry is flat, GTG has two subclasses: Symmetric Teleparallel Equivalent to GR (STEGR) and Teleparallel Equivalent to GR (TEGR).
@@
STEGR is, on one hand, obtained by the imposition of vanishing torsion that is realized in the use of the so-called Stükelberg fields [15]. This theory describes gravity in terms only of the non-metricity. Furthermore, the imposition of the so-called coincident gauge leads to a specific subclass that is equivalent to GR with excepting the second-order derivative terms of its action functional [21, 20]. It is nothing but what Einstein first treated for deriving his field equations when applying the variational principle [22], although it implies also that this subclass of GTG is deprived of the property of the covariance [23]. On the other hand, TEGR is obtained by the imposition of vanishing non-metricity and describes gravity in terms only of the torsion. However, differing from STEGR, unveiling the true identity of vanishing non-metricity from the viewpoint of its gauge structure is still under development. One of the reasons for this situation would be the lack of a unified-description of non-metricity with curvature and torsion. Here, the jargon the “unified-description”, this is nothing but the main subject of the current paper, means that both curvature and non-metricity can be derived from Cartan’s structure equation by using a common connection 1-form. In fact, geometries that are characterized by both curvature and torsion have the unified-description labeled as Riemann-Cartan geometry [24, 25, 26]. In this geometry, these geometric quantities are constructed by the so-called Cartan connection, which is the extension of the Ehresmann connection by using the Möbius representation of a principal bundle with the affine gauge group [27, 28, 14]. This approach to constructing the geometries makes it possible to unveil its gauge structure and the true identity of vanishing torsion (and also vanishing curvature) [20, 29, 30, 31]. Therefore, a unified-description in the same manner that takes also non-metricity into account would open the way for unveiling the true identity of vanishing non-metricity. However, to author’s knowledge, there is no previous work for the unified-description on curvature and non-metricity. In addition to this point, once this approach is realized, the unified-description would make it possible to scrutinize the gauge structure not only of GTG but also of MAG. Furthermore, just alternating the gauge group, the unified-description allows us to easily switch the metric-affine geometry to another geometry. Combining the so-called Inönü-Wigner group contraction [32, 33], the relationships among each geometry could be unveiled, and a new geometry, or equivalently, a new geometric quantity, would also be discovered. The purpose of the current paper is to provide the mathematical fundamentals of a unified-description of curvature, torsion, and non-metricity for the sake of the above investigations.